Louisiana State University Health Sciences Center Administration & Finance
 

From: Ronnie Smith, Vice Chancellor, Administration and Finance
Subject: Hurricane Katrina Recovery – Update #23
Posted: August 18, 2006

The majority of the work to prepare the MEB for the return of students has been completed! The enhancements to the MEB 3rd Floor seminar rooms are complete. The damaged MDL #6 has reopened. Lecture Room B was put back into service on August 10th. The Morgue, the Gross Anatomy Lab and the Swing Lab were ready for use on August 16th. Some punch list items remain to be completed, but these will be addressed during times when the rooms are not in use. The details on all of the repairs within the MEB can be found in specific highlights provided below.

The contractors making the temporary repairs to the Clinic and Administration buildings on the Dental School campus continue to make excellent progress. Punch lists are being developed and provided to the two contractors routinely. As repaired systems are being restarted, some additional damage is being discovered. The contractors are diagnosing the problems as they occur and are proposing the specific additional repair work needed. FEMA and the Office of Facility Planning & Control are working with us to review and approve the additional scope of work, so that work can proceed quickly.

The environmental remediation work at 1542 Tulane is completed, but recent test results show that some limited re-cleaning is still needed. On Monday, August 14th, the LSUHSC faculty and staff were once again allowed to access spaces within 1542 Tulane via the 2nd floor stairwell, by following the normal clearance procedures. Access to the 1st floor and basement is still prohibited until the necessary cleaning is completed and the final clearances are received.

The contractor making temporary repairs to the building’s electrical and mechanical systems at 1542 Tulane has begun receiving critical parts and equipment and the installation of those items has begun. Much of the preparatory work needed for the completion of the planned repairs has been completed.

Some of the first designs for permanent building repairs that include plans for mitigation are being completed and presented to FEMA for their initial consideration. In particular, the plans for the Lion’s/LSU Clinic building were recently shared with FEMA. These plans include the relocation of the Isidore Cohn, M.D. Learning Center from the 1st floor to the 6th floor, and the elevation of critical building components to the 2nd floor where they would be protected from flooding. We have also proposed some protective mitigation measures for the multiple building elevators across the downtown and Dental School campuses.

Here are highlights of work that has been accomplished since our last update was posted two weeks ago.

  • The contractor for the permanent repairs to the MEB Lecture Rooms, the Gross Anatomy Lab, the Morgue, and the Multidisciplinary Labs continues to push towards completion of the contract.
    • The air handling unit supplying the 2nd floor lobby in the MEB is operational. The reheat valves still need to be piped and insulated.
    • Lecture Room B:
      • The room was cleaned and dusted.
      • The A/V equipment has been tested and is operating.
      • The network connections to the individual table top locations will be completed early next week.
    • Lecture Room A:
      • The electricians continued to make the individual electrical and network connections to each of the new table tops in Lecture Room A. Problems with power connections to 14 of the table tops have been resolved.
      • The remaining acoustic wall paper and the trim has been installed.
      • The podium still needs to be connected to the network.
      • The majority of the modesty panels and table tops have been installed.
    • The case work, counter tops and plumbing fixtures have all been installed in the Morgue, the Gross Anatomy Lab and the Swing Lab.
      • All of the sinks in these rooms are operational but one. Repairs to the damaged glass drains will be completed shortly.
      • The rubber base molding must still be installed around the bases of all the new casework.
  • The temporary repairs to the School of Dentistry Clinic and Administration buildings continue.
    • The contractor at the Dental School has taped and floated the drywall for the walls of the new switchgear room.
    • One medical air compressor is ready to operate.
    • The manufacturer of the temporary chiller is planning to have a training class this week on its operation. This will allow them to turn over operation of the chiller.
    • The temporary chiller at the Dental School was started up on 8/8 so that the existing cooling tower could be tested.
      • The east side of the cooling tower is operating with only minimal repairs required.
      • Problems have developed in the testing and operation of the newly installed temporary chiller at the Dental School. The contractor is trouble-shooting the condenser water pumps and the chilled water pumps to determine what additional repairs may be needed. Chiller programming and balancing has been delayed until the condenser pumps & chilled water pump problems are corrected.
    • The contractor is trouble-shooting the loading dock ramp flood gate to determine what needs to be repaired in order to make it operational.
    • The architects and engineers are working on the design for the repairs to the adjacent flood walls.
    • Testing of the Compressed Air System is ongoing – they are continuing to find air leaks at the dental chairs in the Clinic Building.
    • Testing of the buildings’ steam lines are ongoing - a leak was found in one of the pipes in the basement of the Clinic Building.
    • Ventilation fans in the Power Plant were shut-down due to a failure of one of the fans - either the switch or the motor is believed to be at fault.
    • The start-up and testing of the Ice Water Machines/System occurred on 8/15/06.
      • The installation of pipe insulation for the ice water system is complete.
    • Temporary electrical power has been restored to basement.
    • The connection of the wet-vac piping to the building plumbing system is complete. A light fixture that impeded the connection has been relocated. The wet-vac system electrical controls still need to be connected to building power.
    • The hot water system has been connected to the expansion tank.
    • All of the water piping (including steam lines) in the mechanical room of the Clinic Building basement has been insulated. The pipes must still be properly labeled.
    • The removal and destruction of flood damaged medical records that have been held in the Administration building’s basement has begun. Approximately 50% of the records have been removed.
    • The insulation of the water pipes in the Administration building basement is underway and is progressing toward completion. The pipes will still require proper labeling.
    • The original building switchgear has been completely dismantled and removed.
    • Natural gas piping is complete, but has been shut off until the system has been tested. The medical gas piping (NO2, O2, etc.) is progressing well.
    • Mold remediation has begun in some of the upper rooms that were heavily impacted; Rooms 6413B, 6409, & 7409A.
    • The piping for repairing the medical gas system has been delivered and is staged onsite.
    • Replacement of governor cables for Elevator #4 in the Clinic Bldg underway. Elevator doors for Elevator #4 on the 1st floor has been removed...awaiting new replacement doors.
    • The recently emergency chiller failed on 8/18 during brief afternoon thunderstorm. The 24 hour response team was called to bring the chiller back online.
    • Elevator repairs in the clinic building are ongoing. Contractor is working to have the freight elevator and at least one passenger elevator available as soon as possible.
      • Elevator car #3 in the Clinic building has been completely removed.
      • Elevator doors #1, #2, & #4 have been completely removed in the Clinic basement. The contractor is awaiting the arrival of the replacement doors.
      • Elevator travel cables for the elevators in the Clinic building have arrived and are staged onsite.
      • Elevator buffers and travel cables for the elevators in the Administration building are also onsite.
  • Work on the 3rd environmental remediation contract addressing the balance of the buildings and facilities on the downtown campus is substantially complete. Specific updates on progress are as follows:
    • All spaces on the 3rd through the 8th floors have been cleared.
    • All of the spaces on the 2nd floor have been cleared.
    • The majority of the spaces on the 1st floor and in the basement have been cleared.
    • The remediation contractor will be required to return to re-clean the few areas that have not been cleared.
    • LSUHSC faculty and staff may once again access spaces within 1542 Tulane via the 2nd floor stairwell, by following the normal clearance procedure. Access to the 1st floor and the basement is still prohibited.
  • Bids have been received, evaluated and a contract has been awarded for additional repairs to the chilled water pump and related mechanical/electrical components in the Central Utility Plant, attached to the Roman St. Garage. These additional repairs will insure that we have sufficient chilled water capacity for the hottest summer months and also have a redundant pumping system should one fail or require repairs.
    • The variable speed drive and filters are on order. The necessary valves are ready to be replaced.
  • The contractor responsible for the temporary repairs to the building’s critical electrical & mechanical systems at 1542 Tulane Ave. has been making good progress.
    • The contractor has traced wires and stripped the beakers out of electrical panels, preparing various electrical panels for reuse.
    • The old building switchgear has been demolished.
    • The two chiller pumps have been rebuilt and reinstalled.
    • The new wiring was delivered.
    • The damaged electrical panels have been gutted and cleaned. They are now ready for the new wiring to be installed and connected.
    • The 500KV and 1000KV switchgears, LP1 panels, starters and cables have arrived and were lowered into the basement.
    • The contractor has mounted several of the electrical panels needed for the repairs.
    • Old wires and pipes that would interfere with the repairs have been removed.
    • The installation of the temporary switch gear units and the transformers began on 8/16. The location for the installation of these temporary units has been changed so that sufficient space and clearances would be available. The design engineers have approved the revised location.
    • The feeders were run to Panels PLL4 and PLL1.
    • Electrical panels PL2 and 3 were mounted on 8/18.
    • The PL1 cables have been run but are not yet terminated. The contractor is in the process of building the cable supports from the ceiling of the basement.
    • Several men are working on tying in the feeder cables from the 1KV transformer to the 3KV switchgear which powers the elevators.
    • The 600 Amp Disconnect is mounted and the feeder cables are terminated to it.
    • Over the coming weekend the troughs will be mounted for the cables that power the sump pumps.
    • The 13.8KV switchgear will not arrive until Monday, 8/21. A crane has been ordered and will be arrive on site next Monday as well. It will be used to lower the switchgear into the basement and lift old parts and debris out.
      • The lot will be closed Monday and Tuesday, August 21st and 22nd. No vehicles will be allowed in the lot during this process and all vehicles must be removed prior to 5:00 AM Monday, August 21st.
  • The contractor making permanent window repairs to the downtown campus is progressing well.
    • 96 window systems have been replaced.
    • Three window systems are still on order.
    • There are three new windows that are included in a proposed change order that will require approximately 8 weeks to be delivered, once the change is approved.
    • The six damaged windows at 1542 Tulane, which were part of the original project, remain to be replaced.
    • An additional change order is being prepared for submission to the architects. This will address leaks in the atrium glass ceiling in the Resource Center. The revised proposal including all new items will be sent to the architect Monday.
  • Bids for a contract to add new lighting to various Student Parking Lots have been received and are being evaluated. We anticipate that the contract documents will be signed and returned within the next week to 10 days. At that time, the work can begin and demolition will start.
  • Elevator #1 in the Resource Center is back in operation, following repairs completed today. (This was an elevator previously certified, but that had been out of service for some time.)
    • The contractor is also working on correcting electrical problems in Elevator #4. Once these problems are resolved, the elevator will be certified for regular use.
    • The contractor is correcting maintenance items so that the Freight Elevator may be certified.
  • In the MEB cafeteria, the contractor is installing the grill hood exhaust fan which required repairs following the storm.

Please continue to follow our progress through our Daily Updates that are also available on the web site.